OpenStudy (anonymous):

What is the derivative of sin (t/4)?

OpenStudy (emogirlatemoooocow):

f(x)=cos(πt/4) chain rule: derivative of f(g(x))=f'(g(x))g'(x) here f(x)=cos(pi*t/4) and g(x)=(pi/4)t so g'(x)=pi/4 f′(x)=−sin(πt/4)(π/4)=−(π/4)sin(πt/4)